OpenStudy (anonymous):

34. Side dishes You choose any 2 of the following side dishes with your dinner: mashed potatoes, Cole slaw, French fries, applesauce, or rice. How many different combinations of side dishes can you choose?

jimthompson5910 (jim_thompson5910):

There are 5 dishes to choose from at first, then there are 5-1 = 4 dishes left to choose from after you've made the first choice. So there are 5*4 = 20 ways to choose 2 dishes. Since order doesn't matter, you're overcounting each combo twice. So you need to divide by 2 to get 20/2 = 10 So there are 10 combinations.